OpenStudy (anonymous):

Which of the following equations is equivalent to 1/2x - 2/3y = 5? x - 2y = 30 4x - 3y = 30 3x - 4y = 30

OpenStudy (zepp):

What would you get if you multiply every term by 6?

OpenStudy (zepp):

@ik_kishan, you should try to help the person instead of giving straight out the answer.

OpenStudy (anonymous):

i gave the answer coz u helped him....

OpenStudy (zepp):

Okay, let me explain; you interrupted my teaching by giving out the answer and I find this extremely disrespectful, you ain't gonna help him if you only give out the answer.

OpenStudy (inkyvoyd):

@ik_kishan , do you realize that you essentially both disrupted the asker, but you also just stole a medal?

OpenStudy (zepp):

Now, by multiplying everybody by 6, you'll get \(\large \frac{1*6}{2x}-\frac{2*6}{3y}=5*6\) Now, 6/2 = 3 12/3 = 4 Hence, \(3x-4y=30\)